OpenStudy (bohotness):

What is the distance between points (10, 7) and (2, 7) on a coordinate plane? units What is the distance between points (10, 7) and (2, 7) on a coordinate plane? units

OpenStudy (ilovecake):

Why is it written twice?

OpenStudy (bohotness):

i didnt it by mistake

OpenStudy (ilovecake):

oh

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle {\rm D} =\sqrt{ (\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2})^2+(\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2})^2 }}\) where \(\Large\color{black}{ \displaystyle {\rm D} }\) is the distance. \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (bohotness):

O.O can you break it down please?

OpenStudy (solomonzelman):

break it down?

OpenStudy (solomonzelman):

well, your points are \(\Large\color{black}{ \displaystyle (\color{green}{10}~,~~\color{blue}{7}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{2}~,~~\color{red}{7}) }\) are your two points.

OpenStudy (solomonzelman):

is that right, no typos?

OpenStudy (solomonzelman):

with these 2 points you don't need the diatnce formula.

OpenStudy (bohotness):

yes

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

you just need to subtract, because you are just finding the HORIZONTAL distance.

OpenStudy (bohotness):

okay :D

OpenStudy (solomonzelman):

alright, when you are done tell me the answer. If you have question(s) ask.

OpenStudy (bohotness):

8/7

OpenStudy (solomonzelman):

what if I told you to find the distance between 2 and 10, what would tell me the answer is?

OpenStudy (bohotness):

think 6 or 8

OpenStudy (solomonzelman):

well, 10-2 is 8, isn't it?

OpenStudy (bohotness):

yes

OpenStudy (solomonzelman):

So for instance if you walked 2 miles, to complete a walk to 10 miles, you would have to walk another 8 miles.

OpenStudy (bohotness):

yes

OpenStudy (solomonzelman):

lets change to meters.

OpenStudy (solomonzelman):

it is all same, just conceptualizing.

OpenStudy (bohotness):

okay

OpenStudy (solomonzelman):

now, all of this is happening on the line y=7.